# A Joint Learning Framework for Bridging Defect Prediction and Interpretation

### v1

@@ IV-B2 Data preprocessing
We leverage normalization and different data sampling techniques to process the imbalanced data. Specifically, normalization is defined as follows:
@@
where $x$ is the sample of a dataset, and ${\rm min}(\cdot),{\rm max}(\cdot)$ are the minimum and maximum operations applied column-wise across the dataset. The features in the dataset exhibit significant differences in value range. Without normalization, features with larger numerical ranges might unfairly influence the model. Normalization ensures that all features are on the same scale, so the model treats them equally in terms of their impact. Moreover, normalization helps the algorithm converge faster and more smoothly. As shown in the Table III, normalizing the data significantly improves the model’s performance compared to using the raw features. We adopt two sampling techniques to the dataset: SMOTE (Synthetic Minority Oversampling Technique) and RUS (Random Undersampling). The technical details of SMOTE and RUS can be found in [49] and [54]. These techniques address class imbalance in defect prediction tasks but may also affect model interpretability. SMOTE balances the dataset by generating synthetic samples, which may subtly alter the distribution of dataset and distort the contribution of feature to the prediction results. In contrast, RUS reduces the dataset size by randomly removing majority-class samples, potentially simplifying the decision process but at the risk of discarding important information, which may affect the stability of the generated explanations.
TABLE III: Average Performance of the Models with and without Data Preprocessing Across All Datasets
F-measure AUC
Raw Data 0.5822 0.6041
Preprocessed Data 0.6704 0.7109
@@ IV-B3 Statistical significance test
To verify the significance of the improvements achieved by our approach, we introduced $Cohen^{\prime}s\ d$ in this section. $Cohen^{\prime}s\ d$ is used to measure the difference between two groups, which can take any value between 0 and infinity. When $Cohen^{\prime}s\ d$ takes values in the ranges $[0,0.2),[0.2,0.8),and\ [0.8,+\infty)$, it indicates that the differences between the two groups are small, medium, and large 444 https://en.wikipedia.org/wiki/Effect_size. The definition of $Cohen^{\prime}s\ d$ is as follows:
@@
TABLE V: Average CDs of LIME, BreakDown and proposed method
LIME BreakDown Joint_SDT
Average 65% 74.5% 100%
Improvement 53.85%↑ 34.23%↑ -
IV-C Experimental results
The experimental results are presented as follows.
IV-C1 The results of RQ1
To evaluate the consistency of the interpreter trained by our framework (denoted as Joint_SDT), LIME and BreakDown, we employed each of them to interpret an instance ten times, and calculate the average CD for each method. As shown in Table V, the average CD-10% of LIME, BreakDown and Joint_SDT are 65%, 74.5%, and 100%. This result suggests that LIME exhibits poor consistency. To further investigate the consistency of LIME, we visualized its outputs in Fig. 6 and calculated the coefficient of variation (CV) for the weights assigned to different metrics across the ten interpretations. The results show that the CV is 42.18%. The high CV suggests that LIME is unable to consistently assign similar weights to the same metric across multiple explanations. We believe this inconsistency is partly due to the sampling technique and the hyperparameters used by LIME.
(a) The first interpretation
(b) The second interpretation
(c) The third interpretation
(d) The fourth interpretation
(e) The fifth interpretation
(f) The sixth interpretation
(g) The seventh interpretation
(h) The eighth interpretation
(i) The ninth interpretation
(j) The tenth interpretation
Furthermore, we conducted a comparison between the interpretations generated by LIME and Joint_SDT in terms of AI and FI. Since BreakDown does not provide an explicit surrogate function like LIME [8], its AI and FI of BreakDown cannot be calculated. As shown in Table IV, across all projects, Joint_SDT achieves an average improvement of 12.38% in AI and 19.45% in FI compared to LIME. The $Cohen^{\prime}s\ d$ results in the table further confirm that Joint_SDT outperforms LIME, with a large effect size in AI and a moderate effect size in FI.
IV-C2 The results of RQ2
@@
TABLE VI: Prediction Performance of Joint_CNN and Base_CNN in Terms of F-measure and AUC
F-measure AUC
Base_CNN Joint_CNN Base_CNN Joint_CNN
JDT 0.6802 0.8544 0.7297 0.8360
LC 0.5686 0.8285 0.6923 0.6862
ML 0.5686 0.7123 0.6897 0.6418
PDE 0.7475 0.7111 0.7039 0.7032
jm1 0.5755 0.8621 0.6240 0.6280
kc1 0.6443 0.7303 0.7545 0.8087
mc1 0.7352 0.8621 0.8082 0.6280
PC1 0.7475 0.8621 0.7039 0.6280
PC2 0.4996 0.8621 0.5000 0.6280
PC3 0.5588 0.7957 0.7634 0.8324
PC4 0.7021 0.7472 0.8169 0.7843
ant-1.7 0.7991 0.7907 0.8276 0.8022
camel-1.6 0.6932 0.7337 0.5992 0.6687
ivy-1.2 0.8273 0.8347 0.6016 0.8594
jedit-4.1 0.5862 0.6946 0.5429 0.8125
log4j-1.0 0.7755 0.6520 0.8500 0.8500
lucene-2.4 0.5623 0.6947 0.5482 0.7643
poi-3.0 0.7156 0.6951 0.7263 0.7414
synapse1.2 0.6616 0.6486 0.8235 0.7516
xerces-1.3 0.7921 0.8407 0.8498 0.7875
Average 0.6720 0.7706 0.7078 0.7421
Improvement 14.67%↑ - 4.85%↑ -
$Cohen^{\prime}s\ d$ 1.1321 - 0.3553 -
TABLE VII: Average Performances of Joint_CNN and Base_CNN in Terms of PPC, PNPC and FOR
Group PPC PNPC FOR
Base_CNN Joint_CNN Base_CNN Joint_CNN Base_CNN Joint_CNN
AEEEM 0.6570 0.7394 0.3430 0.2606 0.2597 0.1662
NASA 0.6811 0.7708 0.3189 0.2292 0.0503 0.0256
PROMISE 0.4974 0.5167 0.5026 0.4833 0.1867 0.1398
Average 0.6118 0.6756 0.3882 0.3244 0.1656 0.1105
Improvement 10.43% - 19.67$\downarrow$ - 49.86$\downarrow$ -
According to Table VI, we observed that Joint_CNN outperforms Base_CNN in terms of both F-measure and AUC. Across all datasets, Joint_CNN achieves an average improvement of 14.67% in F-measure and 4.85% in AUC compared to Base_CNN. Table VII presents the average performance of Base_CNN and Joint_CNN on three groups: PROMISE, NASA, and AEEEM, evaluated using the metrics PPC, PNPC, and FOR. The results in Table VII show that, compared to Base_CNN, Joint_CNN achieves improvements of 10.43% in PPC, 19.67% in PNPC, and 49.86% in FOR. This indicates that Joint_CNN is more accurate in predicting clean modules, more effective in predicting defective instances on average, and better at reducing false negative instances across the target projects. The $Cohen^{\prime}s\ d$ results indicate a large effect size in F-measure and a moderate effect size in AUC, further confirming the superiority of Joint_CNN over Base_CNN. According to Table VIII, joint_CNN outperforms SVM, RF, DBN and AutoSpearman in terms of both F-measure and AUC. Specifically, Joint_CNN achieves average improvements of 42.86%, 30.63%, 61.75% and 2.13% in F-measure compared to SVM, RF, DBN and AutoSpearman, respectively. Similarly, in AUC, Joint_CNN demonstrates average improvements of 10.93%, 15%, 20.75% and 0.97%. The $Cohen^{\prime}s\ d$ results further validate the performance advantages of Joint_CNN. In terms of F-measure, the effect sizes are large, indicating significant improvements over SVM, RF and DBN. While the effect size for AutoSpearman is smaller, Joint_CNN still demonstrates overall better performance. For AUC, the effect sizes vary across different comparisons. Specifically, the effect size for SVM is medium, while the effect sizes for RF and DBN are large, indicating significant improvements. Although the effect size for AutoSpearman is small, Joint_CNN still demonstrates overall robustness.
TABLE VIII: The F-measure and AUC performance of different models on all datasets
@@
We believe there are two reasons related to the improvement: (1) We explicitly integrate the prediction error $L_{pred}(f,D)$ into the loss function of the joint learning, enabling the predictor to capture discriminative information contained in the data during the training process, and subsequently improving the corresponding accuracy. (2) The $L_{int}$ can be regarded as a penalty term of $L$, which effectively encourages the framework to learn more generalizable representations of the data, leading to improved predictive accuracy for the data that is not contained in training dataset.
IV-C3 The results of RQ3
To demonstrate the global interpretability of our approach, we take the PC2 dataset as an example. As presented in Fig. 5, the decision logic can be abstracted as the path from the “DECISION_DENSITY” metric (denoted as the root) to the leaf node (denoted as 1 or 0). For example, one feasible decision path can be presented as DECISION_DENSITY -> MAINTENANCE_SEVERITY -> PERCENT_COMMENTS -> PERCENT_COMMENTS -> 0, which is similar to the form of decision tree’s interpretation. So, we can conclude our approach achieves the global interpretability.
Furthermore, we calculate the CD, FI and AI values for LC dataset to evaluate the reliability of our approach and SP-Lime. As presented in table IX, our method achieves a better performance in all three indicators, indicating our method is more reliable than SP-Lime. The reason for this result can be partially attributed to the random sampling involved in perturbation processes of SP-Lime [55].
TABLE IX: The CD, FI and AI values of Joint_SDT and SP-Lime
@@ IV-C3 The results of RQ3
IV-C4 The results of RQ4
We conducted an ablation experiment on 5,589 instances in the PC2 dataset to analyze the impact of different features on model performance. The interpretations generated by our approach are presented in Fig. 5. In the interpretations of our approach, the metrics that have a greater impact on the prediction results are located closer to the root node.
Meanwhile, the impacts on the F-measure, AUC, and MCC after removing certain metrics are presented in Fig. 7. We can see that the accuracy of Joint_CNN went down significantly across MCC, F-measure and AUC three indicators after removing the “DECISION_DENSITY”. This indicates that this metric is critical for the accuracy of the predictor. This finding is consistent with the interpretation of our approach, as our approach placed it at the root node of SDT, indicating its decisive role in the model’s decision-making process. Thus, the significant decline in performance after removing this feature further corroborates its key role in the model. The importance of this feature is strongly evidenced both by the performance change and its position in the explanation.
To enhance the comprehensiveness of the analysis, the ablation experiment was extended to all datasets. The most important features interpreted by our approach for each dataset, as well as the ranking of performance degradation on the F-measure, AUC, and MCC after removing these features, are presented in Table X. The results demonstrate that the performance degradation is consistent with our explanatory results, further validating that our method effectively captures key features to improve prediction performance and provide accurate model interpretations.
TABLE X: Ranking of Feature Importance Based on SDT Interpretations and Ablation Results
@@ IV-C4 The results of RQ4
Dataset Top Feature Identified by SDT F-measure Rank AUC Rank MCC Rank
JDT WCHU_numberOfMethodsInherited 4 9 5
LC WCHU_cbo 2 1 1
ML ck_oo_numberOfPublicMethods 5 9 5
@@
xerces-1.3 max_cc 2 8 2
(a) F-measure
(b) AUC
(c) MCC
@@ V Threats to validity
While the proposed model demonstrates advancements in both prediction accuracy and interpretability, certain limitations may affect its generalizability and reliability. The details are as follows.
  • Choice of dataset: The experiment was conducted on the publicly available datasets. These datasets may not fully capture the diversity of software systems in real-world scenarios, potentially limiting the robustness of the model.
  • Hyperparameter setting: The model’s performance is related with the selection of hyperparameters for both the predictor and interpreter. Suboptimal tuning may lead to overfitting or underfitting, resulting in less accurate predictions or interpretations.
  • Application scenarios: In this paper, we focus on the within-project defect prediction that the training data and the test data are generated from the same project. However, the performance of our model in the context of the cross-project setting is not evaluated.
@@ VI Conclusion and future works
In this paper, we explore the possibility of designing a defect predictor and its corresponding interpreter collaboratively. Unlike most existing approaches, which treat defect prediction and interpretation as separate tasks, we consider them strongly correlated. Based on this assumption, we introduce a framework aimed at improving the reliability of interpretation while simultaneously enhancing predictive accuracy by utilizing interpretation results. Based on the proposed framework, we conducted extensive empirical evaluations, which revealed the following key findings:
  • Compared to existing interpretation methods such as LIME and BreakDown, the proposed framework significantly improves the consistency and reliability of interpretations, providing robust explanations from a local perspective. This improvement is attributed to the incorporation of the Knowledge Distillation (KD) principle, which enables the accurate transfer of decision-making knowledge from the predictor to the interpreter.
  • The framework demonstrates superior performance on widely used datasets in terms of both F-measure and AUC, outperforming baseline models such as SVM, RF, DBN, and Base_CNN. By explicitly incorporating interpretation results into the loss function, the framework captures more discriminative information, leading to more robust predictive outcomes.
  • The proposed method offers global interpretability, providing developers with transparent global insights and decision-making rationales. This capability ensures that the framework not only achieves high accuracy but also delivers comprehensive explanatory support for real-world defect prediction tasks.
  • Ablation analysis further demonstrates that the framework effectively identifies key features that have the greatest impact on prediction outcomes, showcasing its ability to provide reliable explanations that align closely with the actual importance of these features.
Future research will investigate several possibilities to further improve the proposed model, including the following:
  • Expand the research by including a wider array of datasets that accurately reflect real-world software systems. Evaluate the robustness and generalizability of our model accordingly.
  • Incorporate automatic hyperparameter optimization methods into our model to reduce the likelihood of suboptimal tuning.
  • Assess the model’s performance in several defect prediction scenarios, such as the cross-project defect prediction.

### v2

@@ IV-B2 Data preprocessing
It involves normalization and data sampling in this section. Specifically, normalization is defined as follows:
@@
where $x$ is the sample of a dataset, and ${\rm min}(\cdot),{\rm max}(\cdot)$ are the minimum and maximum operations applied column-wise across the dataset. Normalization ensures that all features are on the same scale, and it helps the algorithm converge faster and more smoothly. To address the class imbalance issue, we adopt two sampling techniques: SMOTE (Synthetic Minority Oversampling Technique) [3] and RUS (Random Undersampling) [53].
@@ IV-B3 Statistical significance test
To verify the significance of the improvements, we introduce $Cohen^{\prime}s\ d$ in this section, which takes any value between 0 and infinity. When $Cohen^{\prime}s\ d$ takes values in the ranges $0,0.2),[0.2,0.8),and\ [0.8,+\infty)$, it indicates that the improvement is small, medium, and large 444 https://en.wikipedia.org/wiki/Effect_size. The definition of $Cohen^{\prime}s\ d$ is as follows:
@@
TABLE III: The CD-10%s of LIME, BreakDown and proposed method
LIME BreakDown Joint_SDT
Average CD-10% 65% 74.5% 100%
Improvement 53.85%↑ 34.23%↑ -
@@
TABLE V: Prediction Performance of Joint_CNN and Base_CNN in Terms of F-measure, AUC, PPC, PNPC and FOR
F-measure AUC PPC PNPC FOR
Base_CNN Joint_CNN Base_CNN Joint_CNN Base_CNN Joint_CNN Base_CNN Joint_CNN Base_CNN Joint_CNN
JDT 0.6802 0.8544 0.7297 0.8360 0.6000 0.6900 0.4000 0.3100 0.2683 0.1585
LC 0.5686 0.8285 0.6923 0.6862 0.7714 0.7857 0.2286 0.2143 0.1692 0.1538
ML 0.5686 0.7123 0.6897 0.6418 0.6203 0.6952 0.3797 0.3048 0.3216 0.2398
PDE 0.7475 0.7111 0.7039 0.7032 0.7000 0.7867 0.3000 0.2133 0.2105 0.1128
jm1 0.5755 0.8621 0.6240 0.6280 0.5417 0.8036 0.4583 0.1964 0.1335 0.0012
kc1 0.6443 0.7303 0.7545 0.8087 0.5024 0.5308 0.4976 0.4692 0.0783 0.0345
mc1 0.7352 0.8621 0.8082 0.6280 0.8786 0.7818 0.1214 0.2182 0.0012 0.0652
PC1 0.7475 0.8621 0.7039 0.6280 0.6909 0.8182 0.3091 0.1818 0.0500 0.0816
PC2 0.4996 0.8621 0.5000 0.6280 0.9911 0.9875 0.0089 0.0125 0.0018 0.0018
PC3 0.5588 0.7957 0.7634 0.8324 0.7707 0.6752 0.2293 0.3248 0.0397 0.0093
PC4 0.7021 0.7472 0.8169 0.7843 0.8082 0.6644 0.1918 0.3356 0.0167 0.0300
ant-1.7 0.7991 0.7907 0.8276 0.8022 0.7733 0.5733 0.2267 0.4267 0.2267 0.0444
camel-1.6 0.6932 0.7337 0.5992 0.6687 0.5567 0.5876 0.4433 0.4124 0.1692 0.1739
ivy-1.2 0.8273 0.8347 0.6016 0.8594 0.7778 0.7500 0.2222 0.2500 0.0667 0.1000
jedit-4.1 0.5862 0.6946 0.5429 0.8125 0.5312 0.5938 0.4688 0.4062 0.0556 0.0500
log4j-1.0 0.7755 0.6520 0.8500 0.8500 0.5714 0.3571 0.4286 0.6429 0.2727 0.0235
lucene-2.4 0.5623 0.6947 0.5482 0.7643 0.2941 0.3529 0.7059 0.6471 0.4737 0.3684
poi-3.0 0.7156 0.6951 0.7263 0.7414 0.4000 0.5333 0.6000 0.4667 0.3793 0.1724
synapse1.2 0.6616 0.6486 0.8235 0.7516 0.3077 0.6154 0.6923 0.3846 0.2000 0.1111
xerces-1.3 0.7921 0.8407 0.8498 0.7875 0.6304 0.6957 0.3696 0.3043 0.1212 0.1111
Average 0.6720 0.7706 0.7078 0.7421 0.6359 0.6639 0.3641 0.3361 0.1628 0.1022
Improvement 14.67%↑ - 4.85%↑ - 4.4%↑ - 8.33% ↑ - 59.30%↑ -
$Cohen^{\prime}s\ d$ 1.1321 - 0.3553 - 0.0763 - 0.1665 - 0.5338 -
TABLE VI: The F-measure and AUC performance of different models on all datasets
@@
TABLE VII: The CD, FI and AI values of Joint_SDT and SP-Lime
@@
IV-C Experimental results
The experimental results are presented as follows.
IV-C1 The results of RQ1
To evaluate the consistency of Joint_SDT, LIME and BreakDown, we applied each of them to interpret the same instance ten times. As shown in Table III, the average CD-10% scores of LIME, BreakDown and Joint_SDT are 65%, 74.5%, and 100%. This result suggests that both LIME and BreakDown exhibit poor consistency. To further investigate the root cause of such poor consistency, we visualized the interpretations of LIME in Fig. 6. We can see that LIME is unable to assign consistent weights to the same metric across multiple explanations, which we believe it is the root of poor consistency. Moreover, we conducted a comparison between the interpretations generated by LIME and Joint_SDT in terms of AI and FI. Since BreakDown does not provide an explicit surrogate function like LIME [12], we are unable to calculate its AI and FI. As shown in Table IV, Joint_SDT achieves an average improvement of 12.38% in AI and 19.45% in FI compared to LIME. Furthermore, the $Cohen^{\prime}s\ d$ results indicate that the significance of improvement in AI is medium, while in FI, it is large. This results further confirm that Joint_SDT outperforms LIME in the consistency of interpretation.
Although Joint_SDT generally shows improvements in AI and FI, its performance is worse than LIME on certain datasets, such as Log4j. This is largely due to the limited amount of data, where sparse distributions enable LIME to maintain consistent interpretability by reducing sampling randomness. However, the neural network-like structure of Joint_SDT requires a larger volume of data for stable and reliable feature importance estimation. As a result, Joint_SDT struggles to achieve desirable performance on small datasets.
IV-C2 The results of RQ2
According to Table V, we observe that Joint_CNN outperforms Base_CNN in terms of F-measure, AUC, PPC, PNPC, and FOR on most datasets, and the average improvements are 14.67%, 4.85%, 4.4%, 8.33% and 59.30% respectively. The corresponding $Cohen^{\prime}s\ d$ test results indicate that the effect sizes of the improvements in the five indicators are large, medium, small, small, and medium. From these results, we can draw the following conclusions: (1) The improvement of Joint_CNN over Base_CNN across all evaluation metrics indicates that the joint-learning framework makes a positive contribution in improving the performance of SDP models. (2) The improvements in PNPC and FOR suggest that Joint_CNN is more reliable in distinguishing defective and non-defective instances, which is crucial for real-world SDP applications. (3) The constant improvements across all datasets suggests that the joint-learning framework generalize well and is not limited to specific dataset.
Furthermore, according to Table VI, joint_CNN outperforms SVM, RF, DBN and AutoSpearman in both F-measure and AUC. Specifically, Joint_CNN achieves average improvements of 42.86%, 30.63%, 61.75% and 2.13% in F-measure and 10.93%, 15%, 20.75% and 0.97% in AUC. The $Cohen^{\prime}s\ d$ test results indicate that the effect sizes of of improvements in F-measure are large, large, large, and small, while those in AUC are medium, large, large, and small. According to these results, we can confirm that Joint_CNN achieves better accuracy and demonstrates strong robustness across all dataset.
We believe there are two reasons related to the improvements: (1) We explicitly integrate the prediction error $L_{pred}(f,D)$ into the loss function of the joint-learning framework, guiding the predictor to capture discriminative information contained in the data. (2) The revised distillation loss $L_{int}$ can be regarded as a penalty term of $L$, which effectively encourages the framework to learn more generalized features of the data, improving the model’s robustness.
IV-C3 The results of RQ3
To demonstrate the global interpretability of our approach, we use the PC1 dataset as an example. As presented in Fig. 5, the decision logic can be abstracted as the path from the “DECISION_DENSITY” metric (denoted as the root) to the leaf node (denoted as 1 or 0). For example, one feasible decision path can be presented as DECISION_DENSITY -> MAINTENANCE_SEVERITY -> PERCENT_COMMENTS -> PERCENT_COMMENTS -> 0, which is similar to the interpretation format of a decision tree. Furthermore, following the framework of [46], we visualized the AUCs of Joint_CNNs using different metric subsets identified by Joint_SDT and SP-LIME in Fig. 7. The results demonstrate that our method achieves higher AUC values. This confirms that the metric subsets selected by Joint_SDT are more effective for explaining the model’s global decision-making logic compared to those selected by SP-LIME.
Furthermore, we calculate the CD, FI and AI values for LC dataset to evaluate the reliability of our approach and SP-Lime. As presented in table VII, our method achieves a better performance in all three indicators. The reason for this result can be partially attributed to the random sampling involved in perturbation processes of SP-Lime [45].
Figure 7: Comparison of AUC values of Joint_CNNs using metric subsets selected by Joint_SDT and SP-LIME on LC, PC1, and Lucene datasets.
(a) F-measure
(b) AUC
(c) MCC
TABLE VIII: Rankings of Feature Importance
@@ IV-C3 The results of RQ3
Dataset The Most Important Feature Identified by Joint_SDT F-measure Rank AUC Rank MCC Rank
JDT WCHU_numberOfMethodsInherited 4 3 5
LC WCHU_cbo 2 1 2
ML ck_oo_numberOfPublicMethods 3 2 5
@@
xerces-1.3 max_cc 2 5 2
IV-C4 The results of RQ4
According to the structure of Joint_SDT, the most important metric is positioned at the root of the SDT. For instance, in the PC2 dataset, ”DECISION_DENSITY” is identified as the most important metric. To validate this finding, we conduct an ablation experiment that removes each metric selected by Joint_SDT individually and the corresponding changes are observed in the performance of Joint_CNN. The ranking of performance degradation reflects the importance of each metric to the model. The ablation experiment results for PC2 are visualized in Fig.  8. We can see that removing ”DECISION_DENSITY” results in the most significant performance decline across the three indicators: MCC, F-measure, and AUC. This confirms that ”DECISION_DENSITY” is the most important metric for PC2 dataset. This aligns with the interpretation of our method.
To enhance the comprehensiveness of the analysis, we extended the ablation experiment to all datasets. Table VIII presents the most important metrics identified by Joint_SDT across the twenty dataset. The results show that the root metrics identified by Joint_SDT typically correspond to the most significant performance degradations in the ablation experiments.
@@ V Threats to validity
While the proposed model demonstrates improvements in both prediction accuracy and interpretability, certain limitations may affect its generalizability. The details are as follows.
  • Choice of dataset: The experiments were conducted on the publicly available datasets. These datasets may not fully capture the diversity of software systems in real-world scenarios, potentially limiting the generalizability of the model.
  • Hyperparameter setting: The model’s performance is related with the setting of hyperparameters for both the predictor and interpreter. Suboptimal setting may lead to less accurate predictions or interpretations.
  • Application scenarios: In this paper, we focus on the within-project SDP. However, the performance of our model in the context of the cross-project setting has not been evaluated.
@@ VI Conclusion and future works
In this paper, we explore the possibility of designing a defect predictor and its corresponding interpreter collaboratively. Unlike most existing approaches, which treat defect prediction and interpretation as separate tasks, we regard them as strongly correlated. We introduce a framework aiming at improving the reliability of interpretation and enhancing predictive accuracy simultaneously. Based on the extensive empirical evaluations, we obtained the following key findings:
  • Compared to existing interpretation methods such as LIME and BreakDown, the proposed framework significantly improves the reliability of interpretations through the incorporation of the KD principle.
  • The joint-learning framework demonstrates superior accuracy on widely used datasets by explicitly incorporating interpretation results into the loss function.
Future research will explore the following directions:
  • Using diverse real-world datasets to evaluate the robustness and generalization of our method.
  • Integrating automatic hyperparameter optimization to reduce manual tuning efforts.
  • Assess the model’s performance in different defect prediction scenarios, such as cross-project defect prediction.